OpenStudy (anonymous):

Mitch is buying candy bars for his friends. He wants to give 2 bars to each frined, and he wants to have 10 spare bar. he can afford to buy 20 candy bars. What is the algebraic sentence that will help him figure out how man firends he can treat? Let f=the number of friends ?

OpenStudy (anonymous):

my guess was 2f+10=20 correct ?

OpenStudy (anonymous):

Mitch can afford 20 candy so, 20-10 (10 for himself) = 10 to give to friends 10/2 = 5 (# of friends he can give 2 bars) So he can share the candy with 5 friends :)

OpenStudy (anonymous):

you didn't include the variable "f" ..

OpenStudy (anonymous):

@agent0smith